# CodeTaxo: Enhancing Taxonomy Expansion with Limited Examples via Code Language Prompts

### v1

@@ 4.1.1. Datasets
We evaluate the performance of taxonomy expansion methods on small-scale taxonomies using WordNet Sub-taxonomies from (Bansal et al., 2014), and Graphine taxonomies from (Liu et al., 2021a). Specifically, we use 35 Graphine taxonomies with fewer than 100 entities, selected from a total of 227 taxonomies. For the Graphine dataset, we selected 35 taxonomies with fewer than 100 entities out of 227 total taxonomies. In our experiment with WordNet, we utilized 114 sub-taxonomies from the test sets. Additionally, we evaluate three large-scale taxonomies from SemEval-2016 (Bordea et al., 2016) across science, environment, and food domains. Table 1 presents the statistics of these taxonomies, all of which contain entities and definitions curated by human experts. For all benchmarks, 20% of leaf entities are reserved for testing, with the remaining entities used for training.
4.1.2. Baseline Methods.
In the taxonomy expansion task, we evaluated the performance of CodeTaxo against several baseline methods, using both GPT-4o and GPT-4o-mini for in-context learning. The baselines included:
  • TaxoExpan (Shen et al., 2020): adopts GNNs to encode local ego-graphs in taxonomy to enhance entity representation.
  • STEAM (Yu et al., 2020): utilizes the mini-path information to capture the global structure of the taxonomy.
  • HEF (Wang et al., 2022): represents taxonomies as ego-trees to capture hierarchy, fully leveraging the hierarchical structure to improve taxonomy coherence.
  • Musubu (Takeoka et al., 2021): leverages pre-trained models and fine-tunes them as sentence classifiers using queries generated from Hearst patterns.
  • TEMP (Liu et al., 2021b): utilizes a pre-trained model to encode text descriptions of each concept in the taxonomy. It incorporates taxonomic structure information through taxonomy paths.
  • BoxTaxo (Jiang et al., 2023): represent the entities via box embeddings instead of single vector embeddings to capture the hierarchical relation between entities.
  • TaxoPrompt (Xu et al., 2022): adopt prompt tuning on the BERT-based encoder model to capture the taxonomic structure.
  • TaxoInstruct (Shen et al., 2024): a unified framework for taxonomy-related tasks using instruction tuning, focused solely on taxonomy expansion for fair comparison.
4.1.3. Evaluation Metrics.
The performance of CodeTaxo and the baseline models for taxonomy expansion tasks is evaluated using commonly adopted metrics, including accuracy (Acc) and Wu & Palmer similarity (Wu&P), as established in prior work (Yu et al., 2020; Liu et al., 2021b; Wang et al., 2021).
Figure 5. Performance comparison of NL and CodeTaxo (CT) across Llama trained on Code and Natural Language domains. Due to limited contextual window sizes, evaluations were conducted on small-scale sub-taxonomies from WordNet (WN) and Graphine (G).
4.2. Results on Taxonomy Expansion (RQ1)
We compare CodeTaxo with baseline methods for taxonomy expansion, as shown in Table 2. The baselines include two categories: self-supervised methods and in-context learning methods. To our knowledge, CodeTaxo is the first approach to utilize prompting LLMs for taxonomy expansion. We also design a natural language prompt (NL) for comparison to assess the effectiveness of CodeTaxo. The NL prompt mirrors CodeTaxo but omits the entity class definition and uses a sentence for each entity, incorporating the entity’s term, definition, and its parent and child relationships. We have three major observations as follows:
First, both NL and CodeTaxo significantly outperform all self-supervised baselines on the WordNet and Graphine taxonomies. Notably, CodeTaxo in a one-shot setting shows a 72.06% and 103.06% improvement in accuracy over the best self-supervised baselines (TaxoInstruct and TEMP) on WordNet and Graphine, respectively. This indicates that LLM prompting via code prompts can effectively harness the internal knowledge of LLMs, even with limited annotated data. In contrast, self-supervised methods, which heavily depend on annotation data from existing taxonomies, struggle to learn taxonomic relations when the existing taxonomy is insufficiently large for expansion tasks.
Second, for the large-scale taxonomies in SemEval-2016, CodeTaxo in a one-shot setting still outperforms self-supervised methods, even when the existing taxonomy is sufficiently large for generating annotation data. However, the NL prompt does not exhibit the same performance advantages, particularly in terms of accuracy, when compared with self-supervised methods and CodeTaxo. Specifically, CodeTaxo outperforms the best self-supervised baseline, TaxoPrompt, by 10.26%, 8.89%, and 9.21% in accuracy on SemEval-Sci, SemEval-Env, and SemEval-Food, respectively. Conversely, the NL prompt underperforms TaxoPrompt on SemEval-Sci and SemEval-Env, with a decrease in accuracy of 10.75% and 8.53%, respectively. Although NL prompting performs better than TaxoPrompt on SemEval-Food, it still shows a 4.68% decrease in accuracy compared to CodeTaxo. It suggests that CodeTaxo is more effective in representing taxonomic structures than NL and performs better than self-supervised training on existing taxonomies.
Finally, our findings indicate that the performance of CodeTaxo as an in-context learning method is consistently influenced by the inherent capabilities of LLMs and the number of demonstrations included in the prompt. Specifically, we evaluated both NL and CodeTaxo using GPT-4o and GPT-4o-mini. Across all benchmarks, the performance of both NL and CodeTaxo on GPT-4o surpasses that on GPT-4o-mini. Additionally, we assessed 5-shot and 1-shot settings for both NL and CodeTaxo, finding that as the number of high-quality demonstrations increases, the performance of both methods improves across all benchmarks.
4.3. Comparison between NL and CodeTaxo on LLMs and Code-LLMs (RQ2)
Given that CodeTaxo is a prompting method specifically designed for programming languages, we conducted a comparative analysis of its effectiveness against natural language prompting on both general-purpose LLMs and Code-LLMs. To ensure a fair evaluation, we selected models from the Llama family, specifically LLaMa-3-70B-instruct and CodeLLaMA-70B-instruct. Additionally, we included an assessment of CodeLLaMA-34B-instruct to examine the impact of model size on performance. Due to the inability of these LLMs to process large-scale taxonomies within their contextual window, our evaluation was limited to WordNet and Graphine. The experimental results are presented in Figure 5. We have several findings as follows:
First, for both the WordNet and Graphine taxonomies, CodeTaxo demonstrated superior accuracy and Wu&P scores across all tested Code-LLMs and general-purpose LLMs. This suggests that CodeTaxo is more effective in representing taxonomic structures than natural language prompts, not only for black-box general-purpose LLMs like GPT-4 but also for open-source LLMs. Furthermore, whether the LLMs are specialized in code tasks does not appear to affect CodeTaxo’s efficiency, indicating its robustness as a more efficient method compared to natural language prompting.
Secondly, our analysis reveals that CodeTaxo derives greater benefit from Code-LLMs compared to natural language prompts. Specifically, when transitioning from natural language prompting to code language prompting on LLaMa-3-70B-instruct and CodeLLaMA-70B-instruct, CodeTaxo exhibited a 13.33% improvement in accuracy on WordNet, compared to a 6.51% improvement with natural language prompts. Also, we find that the performance of prompting CodeLLaMA-34B-instruct via code language is better on both WordNet and Graphine taxonomy, even though CodeLLaMA-34B-instruct is much more light than LLaMa-3-70B-instruct.
Figure 6. Effect of Top-K relevant entities selected through SimCSE-based Semantic Similarity Filter.
@@ 4.3. Comparison between NL and CodeTaxo on LLMs and Code-LLMs (RQ2)
Table 3. Impact of Entity Definition Sentences (Def.) on CodeTaxo and NL Performance in 1-Shot and 5-Shot Settings.
Table 4. Comparison of average tokens used by NEthical ConsiderationsL and CodeTaxo across 5 benchmarks in 1-shot and 5-shot settings.
4.4. Efficiency Analysis of CodeTaxo (RQ3)
4.4.1. Semantic Similarity Filter Efficiency
This section examines the effect of selecting Top-K entities using the Semantic Similarity Filter on model performance. We conduct experiments using GPT-4o-mini across the three taxonomies in SemEval2016. As shown in Figure 6, across all three taxonomies, increasing the number of Top-K entities generally enhances performance. This improvement is attributed to retaining more entities, which lowers the chance of filtering out the ground truth, thereby boosting prediction accuracy. However, there is a trade-off between search space size and coverage. A smaller search space can improve accuracy by narrowing the model’s focus but risks excluding the ground truth. For example, in the SemEval-Sci, the model achieved optimal performance with a Hit@25 score of 78% by retaining the top 25 entities, demonstrating the filter’s effectiveness in balancing search space and coverage.
In our experiments, we retained the top 50% of entities, ensuring that Hit@n exceeded 90% across all benchmarks. This approach reduces the search space while maintaining a high likelihood of including the ground truth, thereby enhancing the model’s precision in taxonomy expansion.
4.4.2. Number of Tokens
@@
Table 5. Ablation Study of two major modules in the CodeTaxo: All metrics are presented in percentages (%). Configurations indicate whether Demonstration Selection (Demo.) and Semantic Similarity Filter (Filter) were employed.
Query NL Output CodeTaxo Output CodeTaxo Explanation
@@
Def. functional genomics is a field of molecular biology that attempts to describe gene functions and inter-
actions
@@
Exp. Functional genomics is a subfield of genomics that focuses on understanding the functions and interactions of genes within the genome.
@@
Table 6. Case study comparing the outputs of the CodeTaxo and NL prompt using the SemEval-Sci benchmarks. The table presents the definitions (Def.) of each model’s prediction, additionally with the CodeTaxo explanations (Exp.) provided in the last column. Yellow highlights emphasize the specific focus of the query within its definition, as correctly captured by CodeTaxo, while green highlights indicate broader, less precise concepts used by the NL model.
4.5. Ablation Study (RQ3)
4.5.1. Insight of Definition Sentences
@@
4.5.2. Effevtiveness of Demo. Selection and Semantic Similarity Filter
@@
We performed an ablation study on the three SemEval2016 benchmarks mentioned above to assess the effectiveness of the two primary modules in CodeTaxo: Demonstration Selection (Demo.) and the Semantic Similarity Filter (Filter). Due to the relatively small size of the taxonomies in WordNet and Graphine, filtering redundant entities from the existing taxonomies was unnecessary. The results, presented in Table 5, indicate that selecting demonstrations related to the query entity and filtering out unrelated entities in the existing taxonomy significantly improves taxonomy expansion. This finding suggests that incorporating more relevant contextual information and reducing redundant information to narrow the search space is beneficial for both accuracy and the Wu & Palmer (Wu&P) score across all SemEval2016 benchmarks.
4.6. Case Study
This section presents a case study to illustrate the effectiveness of our CodeTaxo framework. We compare the outputs of CodeTaxo and the NL prompt, with model predictions and the corresponding definition detailed in Table 6, where CodeTaxo’s predictions align with the ground truth. We additionally provide the model output using the prompt mentioned in Section 3.2.4, to allow the model to generate explanations and facilitate better discussion.
In the first case, the query functional genomics, which focuses on gene functions and interactions, is correctly classified by CodeTaxo under genomics. The explanation generated by CodeTaxo emphasizes the specific focus of functional genomics within genomics by noting its attention to the “functions and interactions of genes within the genome”, reflecting its specific focus within the broader field. In contrast, the NL model incorrectly selects genetics, a more general term, as the parent entity.
In the second case, CodeTaxo accurately identifies geopolitics as the parent entity of geostrategy, and the explanation provided by CodeTaxo shows its emphasis on geographic factors. The NL model, however, selects the broader category of politics, failing to capture the specific geographic considerations inherent to geostrategy.
These cases demonstrate that CodeTaxo leverages definition information more effectively, enabling a nuanced understanding of taxonomy structures and leading to more precise predictions.
5. Related Works
5.1. Taxonomy Expansion
@@
In taxonomy expansion, various approaches have been developed to integrate emerging entities into existing taxonomies. Aly et al. (Aly et al., 2019) and Ma et al. utilized hyperbolic embeddings to capture taxonomic relations, while Jiang et al. (Jiang et al., 2023) employed box embedding instead of single vector embedding to encode taxonomic relations. Manzoor et al. (Manzoor et al., 2020) introduced implicit edge semantics to enhance entity representations. Self-supervised methods, such as Egonet (Shen et al., 2020), mini-path (Yu et al., 2020), and Ego-Tree (Wang et al., 2021), have also been explored to model structural information within taxonomies. To leverage more semantic information from the textural description of entities, Liu et al. (Liu et al., 2021b), Takeoka et al. (Takeoka et al., 2021) and Xu et al. (Xu et al., 2022) fine-tuned BERT-based models to leverage textual descriptions of entities. Shen et al. (Shen et al., 2024) proposed a unified framework combining various taxonomy construction tasks for instruction tuning. To our knowledge, CodeTaxo is the first work to perform taxonomy expansion via prompting LLMs.
5.2. Code-LLMs for Structured Tasks
@@
Recent studies have demonstrated the strong performance of Code-LLMs in complex reasoning tasks, including symbolic reasoning (Madaan et al., 2022; Cheng et al., [n. d.]), event structure prediction (Wang et al., 2023; Chen et al., 2023), mathematical reasoning (Gao et al., 2023), and knowledge graph construction (Li et al., 2023; Bi et al., 2024). These works highlight Code-LLMs’ ability to transform unstructured text into structured representations, enabling advanced reasoning tasks. In this paper, we focus on enhancing Code-LLMs’ ability to comprehend and expand existing taxonomies through emerging query entities.
6. Conclusion
@@
In this paper, we introduce CodeTaxo, a novel approach to taxonomy expansion that leverages code-based prompts to effectively utilize the inherent knowledge within large language models (LLMs). Our method addresses key challenges in traditional taxonomy expansion by reformulating the task as a code completion problem and employing a Semantic Similarity Filtering mechanism to optimize the use of LLMs’ contextual capacity. Extensive experiments on small-scale and large-scale taxonomies demonstrate that CodeTaxo achieves state-of-the-art performance, in both one-shot settings and five-shot settings. We envision CodeTaxo as a powerful framework for integrating emerging entities into existing taxonomies by accurately identifying appropriate parent entities and also providing new insights for leveraging LLMs in structured knowledge tasks.

### v2

@@ 4.1 Experimental Settings
We evaluate taxonomy expansion on small-scale WordNet sub-taxonomies Bansal et al. (2014) and Graphine taxonomies Liu et al. (2021a). Additionally, we evaluate three large-scale taxonomies from SemEval-2016 Bordea et al. (2016) across science, environment, and food domains. For all benchmarks, 20% of leaf entities are reserved for testing, with the remaining entities used for training. See App. A.1 for details.
Baselines.
We evaluate CodeTaxo, using both GPT-4o and GPT-4o-mini, against self-supervised baselines including TaxoExpan Shen et al. (2020), STEAM Yu et al. (2020), HEF Wang et al. (2022), Musubu Takeoka et al. (2021), TEMP Liu et al. (2021b), BoxTaxo Jiang et al. (2023), TaxoPrompt Xu et al. (2022), and TaxoInstruct Shen et al. (2024), and prompting LLMs through natural language. See details in App. A.2.
@@
Table 2: Impact of Definition Sentences (Def.) on CodeTaxo and NL Performance in 1-Shot and 5-Shot Settings.
Evaluation Metrics:
We use two Accuracy (ACC) and Wu & Palmer similarity (Wu&P) to evaluate the performance of CodeTaxo and baselines. See details in App. A.3
Figure 5: Performance comparison of NL and CodeTaxo (CT) across Llama trained on Code and Natural Language domains. Due to limited contextual window sizes, evaluations were conducted on small-scale sub-taxonomies from WordNet (WN) and Graphine (G).
4.2 Experimental Results
4.2.1 Can CodeTaxo expand taxonomy better than other baselines?
We evaluate CodeTaxo against baseline methods for taxonomy expansion in Table 1, including self-supervised and in-context learning approaches. On WordNet and Graphine, both NL and CodeTaxo significantly outperform self-supervised baselines. In one-shot settings, CodeTaxo improves accuracy by 72.06% and 103.06% over the best self-supervised methods, demonstrating that minimal annotated data effectively unlocks LLMs’ internal knowledge, while self-supervised methods struggle with limited-scale taxonomies. On large-scale SemEval-2016 taxonomies, CodeTaxo surpasses the best self-supervised baseline, TaxoPrompt, by 10.26%, 8.89%, and 9.21% on SemEval-Sci, SemEval-Env, and SemEval-Food, respectively. While the NL prompt underperforms TaxoPrompt on SemEval-Sci and SemEval-Env, it exceeds TaxoPrompt on SemEval-Food but still trails CodeTaxo by 4.68%, highlighting CodeTaxo’s superior ability to capture taxonomic structures. Performance depends on LLM capability and demonstration count, with GPT-4o outperforming GPT-4o-mini and more demonstrations improving accuracy across benchmarks, underscoring the value of high-quality demonstrations for taxonomy expansion.
@@ 4.2.1 Can CodeTaxo expand taxonomy better than other baselines?
Table 3: Ablation Study of two major modules in the CodeTaxo: All metrics are presented in percentages (%). Configurations indicate whether Demonstration Selection (Demo.) and Semantic Similarity Filter (Filter) were employed.
4.2.2 How does CodeTaxo perform across different large language models?
We evaluate CodeTaxo, a prompting method specifically designed for programming languages, by comparing its effectiveness against natural language prompting on both general-purpose LLMs and Code-LLMs. We used Llama-family models, including LLaMa-3-70B-instruct, CodeLLaMA-70B-instruct, and the smaller CodeLLaMA-34B-instruct, to evaluate how model size affects performance.  Given the limited contextual capacity of these models, we focused our evaluation on WordNet and Graphine, as shown in Figure 5. The results highlight CodeTaxo’s superior accuracy and Wu&P scores across all tested models, outperforming natural language prompts in representing taxonomic structures for black-box LLMs like GPT-4 and open-source LLMs. The analysis further reveals that CodeTaxo benefits more significantly from Code-LLMs, with a 13.33% accuracy improvement on WordNet compared to 6.51% for natural language prompts when transitioning to code language prompting. Notably, CodeLLaMA-34B-instruct, despite being smaller, showed better performance on WordNet and Graphine, emphasizing CodeTaxo’s efficiency and robustness.
Figure 6: Effect of Top-K relevant entities selected through SimCSE-based Semantic Similarity Filter.
4.3 Hyperparmeter Analysis of CodeTaxo
This section explores the impact of selecting Top-K entities using the Semantic Similarity Filter on model performance, with experiments conducted using GPT-4o-mini across three SemEval taxonomies. As shown in Figure 6, increasing the number of Top-K entities generally enhances performance by retaining more entities, thereby reducing the likelihood of filtering out the ground truth and boosting prediction accuracy. However, this improvement involves a trade-off: a smaller search space sharpens the model’s focus but increases the risk of excluding the ground truth. For instance, in the SemEval-Sci taxonomy, the model achieved optimal performance with a Hit@25 score of 78% by retaining the top 25 entities, demonstrating the filter’s ability to balance search space and coverage. To further refine this balance, we retained the top 50% of entities in our experiments, ensuring that Hit@n exceeded 90% across all benchmarks.
Query NL Output CodeTaxo Output CodeTaxo Exp.
@@ 4.3 Hyperparmeter Analysis of CodeTaxo
Def. functional genomics is a field of molecular biology that attempts to describe gene functions and interactions
@@
Exp. Functional genomics is a subfield of genomics that focuses on understanding the functions and inter- actions of genes within the genome.
@@
Table 4: Case study comparing the outputs of the CodeTaxo and NL prompt using the SemEval-Sci benchmarks. The table presents the definitions (Def.) of each model’s prediction, additionally with the CodeTaxo explanations (Exp.) provided in the last column (CodeTaxo Exp.). Yellow highlights emphasize the specific focus of the query within its definition, as correctly captured by CodeTaxo, while green highlights indicate broader, less precise concepts used by the NL model.
4.4 Ablation Study
4.4.1 Insight of Definition Sentences
@@
4.4.2 Effevtiveness of Demo. Selection and Semantic Similarity Filter
@@
We performed an ablation study on the three SemEval2016 benchmarks mentioned above to assess the effectiveness of the two primary modules in CodeTaxo: Demonstration Selection (Demo.) and the Semantic Similarity Filter (Filter). Due to the relatively small size of the taxonomies in WordNet and Graphine, filtering redundant entities from the existing taxonomies was unnecessary. The results, presented in Table 3, indicate that selecting demonstrations related to the query entity and filtering out unrelated entities in the existing taxonomy significantly improves taxonomy expansion. This finding suggests that incorporating more relevant contextual information and reducing redundant information to narrow the search space is beneficial for both accuracy and the Wu&P score across all SemEval2016 benchmarks.
4.5 Case Study
This section presents a case study demonstrating the effectiveness of our CodeTaxo framework by comparing its outputs to those of the natural language (NL) prompt, alongside model predictions and corresponding definitions in Table 4. Notably, CodeTaxo aligns closely with the ground truth and generates explanations using the prompt from Section 3.2.4 to facilitate insightful discussions. For instance, in the query functional genomics, CodeTaxo accurately classifies it under genomics, emphasizing its focus on the “functions and interactions of genes within the genome”, whereas the NL model incorrectly selects the broader term genetics. Similarly, CodeTaxo identifies geopolitics as the parent entity of geostrategy, highlighting its emphasis on geographic factors, while the NL model selects the more general category of politics. These cases showcase CodeTaxo’s ability to leverage definition information for a comprehensive understanding of taxonomy structures, resulting in more precise predictions.
5 Related Works
5.1 Taxonomy Expansion
@@
In taxonomy expansion, various approaches have been developed to integrate emerging entities into existing taxonomies. Aly et al. Aly et al. (2019) and Ma et al. utilized hyperbolic embeddings to capture taxonomic relations, while Jiang et al. Jiang et al. (2023) and Xu et al. Xu et al. (2024) employed box embedding and fuse embedding instead of single vector embedding to encode taxonomic relations respectively. Manzoor et al. Manzoor et al. (2020) introduced implicit edge semantics to enhance entity representations. Self-supervised methods, such as Egonet Shen et al. (2020), mini-path Yu et al. (2020), and Ego-Tree Wang et al. (2021), have also been explored to model structural information within taxonomies. To leverage more semantic information from the textural description of entities, Liu et al. Liu et al. (2021b), Takeoka et al. Takeoka et al. (2021) and Xu et al. Xu et al. (2022) fine-tuned BERT-based models to leverage textual descriptions of entities. Zhu et al. Zhu et al. (2023) integrates textual and visual semantics to capture the hierarchical relation between entities. Shen et al. Shen et al. (2024) and Moskvoretskii et al. Moskvoretskii et al. (2024) unified framework combining various taxonomy construction tasks for instruction tuning. To our knowledge, CodeTaxo is the first work to perform taxonomy expansion via prompting LLMs.
5.2 Code-LLMs for Structured Tasks
@@
Recent studies have demonstrated the strong performance of Code-LLMs in complex reasoning tasks Yang et al. ; MA et al. , including symbolic reasoning Madaan et al. (2022); Cheng et al. , graph reasoning Cai et al. (2024), event structure prediction Wang et al. (2023); Chen et al. (2023), mathematical reasoning Gao et al. (2023), and knowledge graph construction Li et al. (2023); Bi et al. (2024). These works highlight Code-LLMs’ ability to transform unstructured text into structured representations, enabling advanced reasoning tasks. In this paper, we focus on enhancing Code-LLMs’ ability to comprehend and expand existing taxonomies through emerging query entities.
6 Conclusion
@@
In this paper, we introduce CodeTaxo, a novel approach to taxonomy expansion that leverages code-based prompts to effectively utilize the inherent knowledge within LLMs. Our method addresses key challenges in taxonomy expansion by reformulating the task as a code completion problem and employing a Semantic Similarity Filtering mechanism to optimize the use of LLMs’ contextual capacity. Extensive experiments on small-scale and large-scale taxonomies demonstrate that CodeTaxo achieves state-of-the-art performance in both one-shot settings and five-shot settings. We envision CodeTaxo as a powerful framework for integrating emerging entities into existing taxonomies by accurately identifying appropriate parent entities and also providing new insights for leveraging LLMs in structured knowledge tasks.
Limitations
This study represents an initial effort to utilize LLMs for taxonomy expansion. Our primary objective is to identify an effective in-context learning strategy to leverage the potential of LLMs. We acknowledge that the performance and scalability of CodeTaxo are constrained by the inherent knowledge of LLMs and the limitations of their context window size. While this paper does not address the challenges of expanding LLM knowledge or increasing context window size, we hope that our work will inspire further research in these areas.
Ethics Statement
Our research addresses taxonomy expansion within general knowledge domains, leveraging our proposed method, CodeTaxo, which uses large language models (LLMs) to generate structured knowledge and overcome the limitations of traditional manual taxonomy construction. We exclusively utilize publicly available datasets and benchmarks, avoiding user-generated, private, or sensitive data to ensure compliance with privacy and ethical standards. While our datasets do not engage directly with ethically sensitive content, LLMs inherently carry biases from their pre-training data, which may influence the structure and content of the expanded taxonomies. To address this, we integrate mechanisms for generating explanatory outputs, enabling detailed scrutiny of the model’s reasoning and identifying potential biases. Additionally, we recognize the risks of applying similar methodologies to subjective or sensitive domains, which could lead to misrepresentation or bias. To mitigate such risks, we emphasize collaboration with domain experts and advocate for responsible application of our methodologies across diverse fields, aiming to promote fairness, accuracy, and ethical research practices.
Appendix A Appendix
A.1 Datasets
We evaluate the performance of taxonomy expansion methods on small-scale taxonomies using WordNet Sub-taxonomies from Bansal et al. (2014), and Graphine taxonomies from Liu et al. (2021a). Specifically, we use 35 Graphine taxonomies with fewer than 100 entities, selected from a total of 227 taxonomies. For the Graphine dataset, we selected 35 taxonomies with fewer than 100 entities out of 227 total taxonomies. In our experiment with WordNet, we utilized 114 sub-taxonomies from the test sets. Additionally, we evaluate three large-scale taxonomies from SemEval-2016 Bordea et al. (2016) across science, environment, and food domains. Table 5 presents the statistics of these taxonomies, all of which contain entities and definitions curated by human experts. For all benchmarks, 20% of leaf entities are reserved for testing, with the remaining entities used for training.
#Concepts #Edges Depth License WordNet 20.5 19.5 3.0 WordNet Graphine 48.2 48.2 4.6 None SemEval-Sci 429.0 451.0 8.0 None SemEval-Env 261.0 261.0 6.0 None SemEval-Food 1,486.0 1,576.0 8.0 None
Table 5: Statistics of five taxonomy benchmarks. For WordNet and Graphine, we report the average for taxonomies included in these two benchmarks.
A.2 Baselines
We compare our method with the following baselines for taxonomy expansion, all experiments are implemented in a server with three NVIDIA A6000 GPUs:
  • TaxoExpan Shen et al. (2020): adopts GNNs to encode local ego-graphs in taxonomy to enhance entity representation.
  • STEAM Yu et al. (2020): utilizes the mini-path information to capture the global structure of the taxonomy.
  • HEF Wang et al. (2022): represents taxonomies as ego-trees to capture hierarchy, fully leveraging the hierarchical structure to improve taxonomy coherence.
  • Musubu Takeoka et al. (2021): leverages pre-trained models and fine-tunes them as sentence classifiers using queries generated from Hearst patterns.
  • TEMP Liu et al. (2021b): utilizes a pre-trained model to encode text descriptions of each concept in the taxonomy. It incorporates taxonomic structure information through taxonomy paths.
  • BoxTaxo Jiang et al. (2023): represent the entities via box embeddings instead of single vector embeddings to capture the hierarchical relation between entities.
  • TaxoPrompt Xu et al. (2022): adopt prompt tuning on the BERT-based encoder model to capture the taxonomic structure.
  • TaxoInstruct Shen et al. (2024): a unified framework for taxonomy-related tasks using instruction tuning, focused solely on taxonomy expansion for fair comparison.
To the best of our knowledge, CodeTaxo represents the first work to address taxonomy expansion using an in-context learning approach. To validate the effectiveness of the code language based prompt design, we additionally propose a prompting method based on natural language prompts. The results obtained using the natural language prompt (NL) are presented in Table 1. To ensure that the natural language prompt communicates the same information as code language prompt in CodeTaxo, we represent each entity using natural language to describe its surface name, definition, parent, and children list. The details of the NL prompt are provided in Table 7. For a more direct comparison, we also demonstrate CodeTaxo’s predictions on the same example in Table 8.
A.3 Evaluation Metrics.
The performance of CodeTaxo and the baseline models for taxonomy expansion tasks is evaluated using commonly adopted metrics, including accuracy (Acc) and Wu & Palmer similarity (Wu&P), as established in prior work Yu et al. (2020); Liu et al. (2021b); Wang et al. (2021). Since CodeTaxo is a generation-based method rather than a ranking-based one, the mean reciprocal rank (MRR) used in the baselines is not applicable to CodeTaxo.
Table 6: Comparison of average tokens used by NL and CodeTaxo across 5 benchmarks in 1-shot and 5-shot settings.
A.4 Efficiency Analysis of CodeTaxo
Token Consumption
Natural Language Prompt
User: Given the current taxonomy, find the parent of the query node. Please note that the query node may be a new node not in the current taxonomy. The parent of given query node always exists, so do not generate ’none’ or ’not found’. You only need to answer the entity name and do not generate any additional content or comments.
lunacy: obsolete terms for legal insanity; parent: insanity; children: [].
irrationality: the state of being irrational; lacking powers of understanding; parent: insanity; children: [].
dementia: mental deterioration of organic or functional origin; parent: insanity; children: [’presenile dementia’, ’alcoholic dementia’, ’senile dementia’].
alcoholic dementia: dementia observed during the last stages of severe chronic alcoholism; involves loss of memory for recent events although long term memory is intact; parent: dementia; children: [].
Pick’s disease: a progressive form of presenile dementia found most often in middle-aged and elderly women and characterized by degeneration of the frontal and temporal lobes with loss of intellectual ability and transitory aphasia; parent: presenile dementia; children: [].
derangement: a state of mental disturbance and disorientation; parent: insanity; children: [].
craziness: informal terms for insanity; parent: insanity; children: [].
presenile dementia: dementia with onset before the age of 65; parent: dementia; children: ["Pick’s disease"].
senile dementia: dementia of the aged; results from degeneration of the brain in the absence of cerebrovascular disease; parent: dementia; children: [].
insanity: relatively permanent disorder of the mind; parent: None; children: [’irrationality’, ’dementia’, ’craziness’, ’derangement’, ’lunacy’].
Query node: Alzheimer’s disease
The parent of query node:
Assistant: dementia
Ground Truth: presenile dementia
Table 7: Example of Natural Language (NL) Prompt.
Code Prompt
User: Complete the next line of code according to the comments and the given code snippet. You need to find the parent of the query node in the given current taxonomy and use the add_parent function. The parent of given query node always exists in the given current taxonomy, so do NOT generate node that is NOT in the given current taxonomy. Note that you only need to complete the next ONE line of code, do not generate any additional content or comments.
from typing import List
class Entity: def __init__(self, name: str, description: str, parent: str, child: List[’Entity’]): self.name = name self.description = description self.parent = parent self.child = child def add_parent(self, parent: ’Entity’): self.parent = parent.name parent.add_child(self) def add_child(self, child: ’Entity’): self.child.append(child)
# Creating entities and establishing parent-child relationship lunacy = Entity(name=’lunacy’, description=’obsolete terms for legal insanity’, parent=insanity, child=[]) irrationality = Entity(name=’irrationality’, description=’the state of being irrational; lacking powers of understanding’, parent=insanity, child=[]) dementia = Entity(name=’dementia’, description=’mental deterioration of organic or functional origin’, parent=insanity, child=[’presenile dementia’, ’alcoholic dementia’, ’senile dementia’]) alcoholic_dementia = Entity(name=’alcoholic dementia’, description=’dementia observed during the last stages of severe chronic alcoholism; involves loss of memory for recent events although long term memory is intact’, parent=dementia, child=[]) Pick’s_disease = Entity(name=’Pick’s disease’, description=’a progressive form of presenile dementia found most often in middle-aged and elderly women and characterized by degeneration of the frontal and temporal lobes with loss of intellectual ability and transitory aphasia’, parent=presenile dementia, child=[]) derangement = Entity(name=’derangement’, description=’a state of mental disturbance and disorientation’, parent=insanity, child=[]) craziness = Entity(name=’craziness’, description=’informal terms for insanity’, parent=insanity, child=[]) presenile_dementia = Entity(name=’presenile dementia’, description=’dementia with onset before the age of 65’, parent=dementia, child=["Pick’s disease"]) senile_dementia = Entity(name=’senile dementia’, description=’dementia of the aged; results from degeneration of the brain in the absence of cerebrovascular disease’, parent=dementia, child=[]) insanity = Entity(name=’insanity’, description=’relatively permanent disorder of the mind’, parent=None, child=[’irrationality’, ’dementia’, ’craziness’, ’derangement’, ’lunacy’])
# creating query node Alzheimer’s_disease = Entity(name=’Alzheimer’s disease’, description=’a progressive form of presenile dementia that is similar to senile dementia except that it usually starts in the 40s or 50s; first symptoms are impaired memory which is followed by impaired thought and speech and finally complete helplessness’, parent=None, child=[])
# Finding the parent of query node
Assistant: Alzheimer’s_disease.add_parent(presenile_dementia)
Ground Truth: presenile dementia
Table 8: Example of Code-based Prompt.